OpenStudy (anonymous):

Find all solutions to the equation. cos2x + 2 cos x + 1 = 0

OpenStudy (lyrae):

\[\cos(2x) = \cos^2(x) - \sin^2(x)\]\[1 = \cos^2(x) + \sin^2(x)\]

OpenStudy (anonymous):

I knew you had to use sine but I wasn't sure how or why.. still a little confused

OpenStudy (lyrae):

Try and substitute with those two, simplify and tell me what you get.

OpenStudy (anonymous):

okayy hold on

OpenStudy (lyrae):

The over all goal is to simplify as much as possible so we get an easy expression to solve for x.

OpenStudy (anonymous):

I got a weird simplification for the first equation I don't think I'm doing it right

OpenStudy (lyrae):

And some times you have to substitute with trig-identities to achieve that.

OpenStudy (lyrae):

Okay write it here and we'll see

OpenStudy (lyrae):

Just the expression after substitution without any simplification.

OpenStudy (anonymous):

I feel like it's too far off to even write. I'm not completely sure how I'm supposed to simplify the terms

OpenStudy (lyrae):

cos(2x) + 2 cos(x) + 1 = 0 \[\cos^2(x) - \sin^2(x) + 2 \cos(x) + \cos^2(x) + \sin^2(x) = 0\]The sines cancel and you get \[2\cos^2(x) + 2 \cos(x) = 0\]\[\cos (x) (\cos (x) + 1) = 0\]

OpenStudy (lyrae):

From there you'll find all solutions when cos(x) = 0 or cos(x) + 1 = 0

OpenStudy (anonymous):

So which parts do I need to solve to determine if it equals 0 for cos(x) or cos(x)+1?

OpenStudy (lyrae):

x = arccos(0) or x = arccos(-1).

OpenStudy (lyrae):

Dont forget that both cos(90 deg) and cos(270 deg) is 0.

OpenStudy (anonymous):

arccos(0)= pi/2 arccos(-1)= pi

OpenStudy (anonymous):

So then I plug those into cos(x)=0 and cos(x)+1=0 ?

OpenStudy (lyrae):

That's only necessary if you want to check if the answer is cocrrect. The assignment is to find the value of x, right? And one soulution is x = arccos(0) => x = 90 deg = pi/2 x = 270 deg = 3pi/2 x = arccos(-1) => x = 180 deg = pi

OpenStudy (anonymous):

So there are a total of three solutions which are pi/2, 3pi/2, and pi?

OpenStudy (lyrae):

No, there are actually an infinte amount of soulutions and that's one possible soulution.

OpenStudy (lyrae):

Finding the other ones are simple though.

OpenStudy (lyrae):

cos is periodic. -1 = cos(pi) = cos(3pi) = cos(9pi) and so on so all you have to do is to add a term for each partial soulution to cover for the periodicity. x = arccos(0) => x = pi n - pi/2 for any whole number n x = arccos(-1) => x = 2n pi - pi for any whole number n x = 2k pi + pi for any whole number k